OpenStudy (anonymous):

Which of the following is equivalent to log 8 16 ? Choose one answer. a. 2 b. 1.248 c. 4 over 3 d. 3 over 4

OpenStudy (anonymous):

Do you mean \(\log_8 16\)?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Then you want to find the exponent that eight raise to that exponent equals 16 \( 8^x=16\)

OpenStudy (anonymous):

so would it be b

OpenStudy (anonymous):

Can you eliminate any of the choices?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

nope trying to find one

OpenStudy (anonymous):

k got it 4/3

OpenStudy (anonymous):

Ok, choice b is between 1 and 2, so \(8^x\) would be between 8 and 64.

OpenStudy (anonymous):

as far as i know: \[\log_{x}a=\ln x/\ln a\] !!

OpenStudy (anonymous):

Got the anwer can you help me on another one please?

OpenStudy (anonymous):

Which of the following is equivalent to log 3y 2x ? Choose one answer. a. common log of 2x over common log of 3y b. common log of 3y over common log of 2x c. log2x over 3y d. log3y over 2x

OpenStudy (anonymous):

I can try.

OpenStudy (anonymous):

Is there a minus sign missing?

OpenStudy (anonymous):

nope the 

OpenStudy (anonymous):

OK, 0x58 gave you the formula above.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Six or seven posts ago, 0x58 posted the change of base formula.

OpenStudy (anonymous):

ok